OpenStudy (anonymous):

If sin x = -(sqrt2)/2)) and cos x = (sqrt2)/2, the measure of angle x is? a. 45 b. 135 c. 225 d. 315

OpenStudy (anonymous):

sin x = -(sqrt2)/2)) cos x = (sqrt2)/2 sin = -ve while cos=+ve Therefore x is in quadrant IV tanx = -1 x= 315

OpenStudy (anonymous):

Since sin x is negative and cos x is positive, x is in quadrant 4. Only angle 315 is in quadrant 4 so d is the correct answer.

OpenStudy (anonymous):

Thank you both.